Question 2 Mario’s executor determined that the estate tax liability for Mario’s estate is $600,000. However, Mario’s executor forgot to file the estate tax return and filed and paid 65 days late. Calculate the penalties that Mario’s estate will now have to pay.

A. $81.000.

B. $90,000.

C. $99,000.

D. $690,000.
Answer #1

Option B, $90,000.

Reason: Penalty = 600,000 *0.5% * 3 months = 9000

Adjusted failure to pay = $81,000

Total penalty = 9000 +81000 = 90,000

Any doubt please comment, If it helped you, please give a thumbs up.